# COMAE: COMprehensive Attribute Exploration for Zero-shot Hashing

### v3

@@ 2.3. Pair-wise Attribute Consistency
where $\epsilon$ is a hyper-parameter and is set through multiple experiments. We randomly sample $k$ negative samples for the $i$-th image, as considering all negative samples will bring high complexity. We get the sample set $S_{d}(i)=S^{+}_{d}(i)\cup S^{-}_{d}(i)$. Afterwards, we formulate the attribute-level contrastive embedding objective as follows,
@@ 2.5. Hashing Learning
where $W_{h,w,c}$ is the learnable transform matrix. It changes the last fully connected layer to a $d$-dimension representation.
To compress intra-class hashing code with strong constraints, we use the hypersphere loss (Liu et al., 2017; Wang et al., 2018) to optimize the FCH layer with large angular boundaries between classes. The spherical hashing objective offers several advantages within the context of similarity-preserving hash coding. One notable advantage lies in its ability to enforce constraints on binary codes such that they reside on a hypersphere in the Hamming space. This geometric constraint enhances the discriminative power of the learned hash codes by facilitating a more structured and separable representation of data points. Additionally, the hypersphere constraint promotes semantic similarity preservation, as points closer to the hypersphere correspond to more semantically similar items. This results in improved retrieval performance, particularly in scenarios where maintaining semantic relationships among binary codes is crucial, such as image retrieval and nearest neighbor search. Furthermore, the geometric regularity induced by the hypersphere constraint contributes to robustness against noise and variations, thereby enhancing the overall effectiveness and stability of the hashing model. It is defined as,
(8) $\mathcal{L}_{\text{hash}}$ $=-\frac{1}{N}\sum_{i}^{N}\log\left(\frac{e^{||b_{i}||\psi(\theta_{yi,i})}}{e^{||b_{i}||\psi(\theta_{yi,i})}+\sum_{j\neq y_{i}}e^{||b_{i}||\cos(\theta_{j},i)}}\right),$
$\psi(\theta_{y_{i},i})=(-1)^{k}\cos(m\theta_{y_{i},i})-2k,$
where $m\geq 1$ is an integer that controls the size of the angular margin, and $\theta_{y_{i},i}\in[\frac{k\pi}{m},\frac{(k+1)\pi}{m}],k\in[0,m-1]$. Compared with other hashing methods, it can achieve a larger separation and a smaller variation of inter-classes.
For a comprehensive exploration between the attribute learning and image hashing tasks, the overall objective is formalized as follows,
@@
where $\mathcal{L}_{i}$ denotes the pointwise-, pairwise-, and classwise-losses respectively, and $\lambda_{i}$ denotes hyper-parameters. The parameters of the hashing network could not be updated by the standard back-propagation as the derivation of $\text{sign}(\cdot)$ is indifferentiable at zero. To solve this issue, the $\text{tanh}(\cdot)$ function is adopted to approximate the results of $\text{sign}(\cdot)$. In this manner, we can optimize the original indifferentiable loss by the mini-batch Stochastic Gradient Descent (SGD).
@@
Methods 24 bits 48 bits 64 bits 128 bits 24 bits 48 bits 64 bits 128 bits 24 bits 48 bits 64 bits 128 bits
LSH 0.0106 0.0151 0.0204 0.0306 0.0055 0.0069 0.0076 0.0095 0.0598 0.0738 0.0764 0.0880
SH 0.1833 0.2729 0.2955 0.3441 0.0568 0.081 0.0886 0.1191 0.0738 0.0811 0.0837 0.0880
ITQ 0.1999 0.2821 0.2964 0.3764 0.0533 0.0765 0.0892 0.1182 0.0725 0.0899 0.0918 0.1007
IMH 0.1282 0.1536 0.1613 0.1681 0.0330 0.0361 0.0364 0.0386 0.0572 0.0668 0.0638 0.0651
PCA 0.2165 0.2530 0.2701 0.2719 0.0547 0.0598 0.0632 0.0695 0.0808 0.0868 0.0891 0.0980
HashNet 0.2086 0.2386 0.2516 0.2749 0.0528 0.0566 0.0595 0.0633 0.0746 0.0825 0.0865 0.0944
GreedyHash 0.3420 0.4169 0.4240 0.4639 0.1132 0.1707 0.1841 0.2326 0.0873 0.1259 0.1398 0.1550
JMLH 0.3607 0.4364 0.4408 0.4711 0.1078 0.1555 0.1987 0.2310 0.0860 0.1168 0.1345 0.1500
ADSH 0.3360 0.4787 0.5105 0.5454 0.0858 0.1607 0.1827 0.2424 0.0886 0.1230 0.1341 0.1731
CSQ 0.3194 0.3988 0.3773 0.4072 0.0996 0.1588 0.1712 0.2201 0.0915 0.1211 0.1375 0.1669
DPN 0.1783 0.2086 0.2378 0.2565 0.0445 0.0728 0.0772 0.1003 0.0737 0.0825 0.0894 0.1126
BiHalf 0.3440 0.4036 0.4223 0.4577 0.0794 0.1280 0.1573 0.2142 0.0077 0.0490 0.0512 0.1307
OrthoCos 0.1709 0.2302 0.2312 0.2566 0.0451 0.0660 0.0736 0.0984 0.0728 0.0809 0.0868 0.1126
CIBHash 0.2113 0.2304 0.2481 0.2618 0.0351 0.0404 0.0411 0.0453 0.0690 0.0808 0.0809 0.0883
TBH 0.0941 0.1201 0.1073 0.1730 0.0157 0.0176 0.0226 0.0252 0.0169 0.0230 0.0248 0.0287
TSK 0.2262 0.3109 0.3873 0.4151 0.0739 0.1200 0.1394 0.1112
ASZH 0.2619 0.3787 0.4032 0.4158 0.0764 0.1192 0.1294 0.1727 - - - -
SitNet 0.2344 0.2406 0.2549 0.2650 0.0880 0.1127 0.1141 0.1167 - - - -
OPZH 0.1056 0.1390 0.1618 0.1961 0.0632 0.0879 0.0962 0.1143 - - - -
AH 0.2275 0.1989 0.3154 0.3557 0.0480 0.0897 0.1089 0.1445 - - - -
COMAE 0.3819 0.4792 0.5133 0.5465 0.1136 0.1777 0.1994 0.2519 0.0923 0.1264 0.1415 0.1741
@@ 3. Experiments
  • RQ1: Can the proposed COMAE gain superior performance in comparison to other baselines?
  • RQ2: How does the performance of COMAE evolve in response to the change in the ratio of unseen classes in the training process?
  • RQ3: What insights can be derived from an ablation study conducted on the proposed approach?
  • RQ4: How does the performance of COMAE evolve in response to the change of hyper-parameters?
@@ 3.1. Experimental Settings
Figure 3. mAP@all and AUC scores and their decline percentage with the change of unseen classes ratio in the training process.
3.1.1. Datasets
In our investigations, we meticulously analyzed three distinguished zero-shot hashing datasets referenced in (Shi et al., 2022), i.e., AWA2 (Xian et al., 2018), CUB (Wah et al., 2011), SUN (Patterson and Hays, 2012). AWA2 encompasses 37,322 images across 50 animal categories, with a configuration of 40 seen and 10 unseen classes, each described by 85 attributes [39]. The CUB dataset includes 11,788 images from 200 bird species, categorized into 150 seen and 50 unseen classes, each characterized by 312 attributes. SUN dataset, on the other hand, contains 14,340 images from 717 scene categories, with a division of 645 seen and 72 unseen classes, each depicted by 102 attributes.
3.1.2. Baselines
  • Image Shallow Hashing methods often leverage handcrafted features or employ traditional machine-learning algorithms to generate compact binary representations. Typical works include LSH (Andoni and Indyk, 2008), SH (Weiss et al., 2008), ITQ (Gong et al., 2012), SDH (Shen et al., 2015a), IMH (Shen et al., 2015b), and PCA (Maćkiewicz and Ratajczak, 1993). Shallow methods are computationally efficient and exhibit simplicity in their design, yet they may struggle to capture complex and high-level semantic features inherent in images.
  • Image Deep Hashing methods, leverage deep neural networks to automatically learn hierarchical and abstract representations of images. Representative works can be categorized into two paths, supervised hashing works, i.e, HashNet (Cao et al., 2017), GreedyHash (Su et al., 2018), JMLH (Shen et al., 2019), ADSH (Zhou et al., 2019), CSQ (Yuan et al., 2020), DPN (Fan et al., 2020), OrthoCos (Hoe et al., 2021), and unsupervised hashing works, i.e, BiHalf (Li and van Gemert, 2021), TBH (Qiu et al., 2021), CIBHash (Qiu et al., 2021). By employing end-to-end learning, deep hashing methods aim to optimize hash codes directly from raw pixel values, allowing for the discovery of intricate and discriminative image representations.
  • Zero-shot Image Hashing methods extend the above algorithms to handle scenarios where the model encounters classes during testing that were not seen during training. These methods often incorporate semantic information, such as image attribute descriptions, or meta-learning strategies to enhance the model’s adaptability to unseen classes. Representative works include TSK (Yang et al., 2016), ASZH, SitNet, OPZH, and AH. ASZH (Shi et al., 2022) uses category labels to learn the hash functions from the seen training data, to generate binary codes for unseen data. SitNet (Guo et al., 2017), and OPZH (Zhang et al., 2019) adopt semantic vectors to find semantic embedding space for knowledge transferability. AH (Xu et al., 2017) constructs relationships and image features for transferability.
3.1.3. Evaluation Metric
For evaluating different aspects of our study, we employed a variety of metrics. These include the mean Average Precision (mAP) and the Area Under the Precision-Recall Curve (AUC) for assessing the ranking quality of search outcomes. Specifically, we present findings for the top 5,000 samples and for the entire dataset, denoted as mAP@5000 and mAP@all, respectively. Additionally, we utilized four conventional metrics: mean Average Precision (mAP), Precision-Recall Curve (PR Curve), Precision@N Curve (P@N Curve), and Recall@N Curve (R@N Curve), with a particular focus on the ranking performance for the initial 5,000 retrieved samples.
3.1.4. Implementation Details
All comparative baselines are conducted with their default hyper-parameter configurations. For a fair comparison, their image feature extractors are substituted with the ResNet101 (Khan et al., 2018). The training epoch is set to 10 and the batch size is set to 64, and more than 10 will be overfitting, resulting in a decrease in accuracy. We take Adam as the optimizer weight decay 0.0005 and the learning rate is set to 0.0001. The $\epsilon$ in Eqn. 9 is set to 0.9, the $\lambda_{1},\lambda_{2}$, and $\lambda_{3}$ in Eqn. 4 are set to 10, 1, and 10 respectively. To ensure the statistical significance of our experimental data, the average outcomes of each technique were calculated based on ten iterations.
(a) PR Curve (AWA2)
(b) PR Curve (CUB)
(c) PR Curve (SUN)
(d) P@N Curve (AWA2)
(e) P@N Curve (CUB)
(f) P@N Curve (SUN)
(g) R@N Curve (AWA2)
(h) R@N Curve (CUB)
(i) R@N Curve (SUN)
3.2. RQ1: Overall Performance
@@
Table. 1 displays the results of the COMAE and all competitive baselines on datasets AWA2, CUB, and SUN with hash code lengths of 24, 48, 64, and 128. Note that the source codes of ASZH, SitNet, AH, and OPZH are unavailable, thus we use their results in paper (Shi et al., 2022). The experimental settings of Table 1 are the same as (Shi et al., 2022). In Fig. 4, we illustrate three distinct curves, i.e., PR Curve, R@N Curve, and the P@N Curve, to evaluate the performance of hash codes with a length of 64 bits on all datasets. Typically, the efficacy of a method is gauged by the PR curve approaching the upper right quadrant, with higher positions on the P@N and R@N curves indicative of superior performance. Overall, we find the following conclusions,
  • As shown in Table. 1, deep hashing models perform better than shallow hashing ones in general, which shows the superiority of neural networks. In addition, The zero-shot hashing models exhibit superior performance in comparison to both deep and shallow hashing models.
@@
  • COMAE surpasses all existing zero-shot hashing benchmarks, in terms of all bit lengths. At the same time, the performance of COMAE remains relatively stable when the number of bits increases. Compared with AH (Xu et al., 2017), an attribute-based zero-shot hashing method, we gain better performance, which serves to validate the correctness and effectiveness of the motivation we put forth.
  • An insightful observation can be derived from the graphical representation in Fig. 4 in all datasets. COMAE attains a heightened accuracy at a comparatively lower recall rate, thereby demonstrating the superiority of our proposed method.
3.3. RQ1: Hashing Code Visualization
In this experiment, we adopted a more intuitive approach to visualizing hashing codes to evaluate the effectiveness of our proposed COMAE. Throughout the comparative experiments, we initially transformed the 64-bit hashing codes into lower-dimensional spaces, subsequently projecting these codes into a 2-D plane using the t-SNE (Van der Maaten and Hinton, 2008) tool. This methodology facilitated a clearer observation of COMAE’s performance through visual means. We present the visualization of hash code for all testing samples with unseen classes on the AWA2 dataset. Results are shown in Fig.  7. According to the visualization results, we find that a) HashNet, DPN and OrthoCos perform badly, which are unable to distinguish the unseen classes. b) Although CSQ and BiHalf gain favorable outcomes in visualization and mAP@5000 results in Table. 1, they still encounter challenges such as substantial inter-class overlap and indistinct boundaries (the classes represented by the green and other colors). c) Compared with other baselines, the proposed COMAE achieves the best performance, as the generated hash codes are more distinguished. It indicates that COMAE can effectively classify images.

### v5

@@ 3.3. Pair-wise Attribute Consistency
where $\epsilon$ is a hyperparameter and is set through multiple experiments. We randomly sample $k$ negative samples for the $i$-th image, as considering all negative samples will bring high complexity. We get the sample set $S_{d}(i)=S^{+}_{d}(i)\cup S^{-}_{d}(i)$. Afterwards, we formulate the attribute-level contrastive embedding objective as follows,
@@
Algorithm 1 Algorithm of COMAE
1:Training Images $X=\{x_{i}\}_{i=1}^{N}$, class labels $Y=\{y_{i}\}_{i=1}^{N}$, and image attribute vectors $A=\{a_{i}\}_{i=1}^{N}$.
2:The parameters $\Theta$
3:for $i=0,...,N$ do
4:     Generate image initial embeddings $f(x_{i})$ based on the image encoder, i.e., ResNet101;
5:end for
6:Generate the predicted prototypes $\{M^{k}\}_{k=1}^{K}$ for each attribute.
7:for $i=0,...,N$ do
8:     Construct the positive sample set $S^{+}_{d}(i)$ of the $i$-th image and negative sample set $S^{-}_{d}(i)$, as Eqn. (3).
9:end for
10:while not converge do :
11:     Compute the pointwise loss, as Eqn. (2).
12:     Compute the pairwise loss, as Eqn. (4).
13:     Compute the classwise loss, as Eqn. (5).
14:     for $i=0,...,N$ do
15:         Generate binary hashing codes $B=\{b_{i}\}_{i=1}^{N}$ based on image representations $f(x_{i})$ through Eqn. (6).
16:     end for
17:     Compute the joint loss through Eqn. (7)
18:     Compute the backpropagation error.
19:     Update the parameters $\Theta$ and image hidden representations $f(x_{i})$.
20:end while
21:return Image hashing codes $B=\{b_{i}\}_{i=1}^{N},b_{i}\in\{-1,1\}^{l}$.
@@ 3.5. Hashing Learning
where $W_{h,w,c}$ is the learnable transform matrix. It changes the last fully connected layer to a $d$-dimension representation. To compress intra-class hashing code with strong constraints, we use the hypersphere loss (Liu et al., 2017; Wang et al., 2018) to optimize the FCH layer with large angular boundaries between classes as $\mathcal{L}_{\text{hash}}$. Compared with other hashing methods, it can achieve a larger separation and a smaller variation of inter-classes.
For a comprehensive exploration between the attribute learning and hashing tasks, the overall objective is formalized as follows,
@@
where $\mathcal{L}_{i}$ denotes the pointwise-, pairwise-, classwise-, and hashing losses respectively, and $\lambda_{i}$ denotes hyper-parameters. The parameters of the hashing network could not be updated by the standard back-propagation as the derivation of $\text{sign}(\cdot)$ is indifferentiable at zero. Thus $\text{tanh}(\cdot)$ is adopted to approximate the results of $\text{sign}(\cdot)$. In this manner, we can optimize the original non-differentiable loss by the mini-batch Stochastic Gradient Descent (SGD). The pseudocode of COMAE is shown in the Algorithm. 1.
3.6. Theoretical Analysis
Here we provide a theoretical analysis for our proposed COMAE.
Theorem 1.
[…]
Theorem 2.
[…]
@@ 3.6. Theoretical Analysis
Methods Reference 24 bits 48 bits 64 bits 128 bits 24 bits 48 bits 64 bits 128 bits 24 bits 48 bits 64 bits 128 bits
LSH CACM08 0.0106 0.0151 0.0204 0.0306 0.0055 0.0069 0.0076 0.0095 0.0598 0.0738 0.0764 0.0880
SH NeurIPS08 0.1833 0.2729 0.2955 0.3441 0.0568 0.081 0.0886 0.1191 0.0738 0.0811 0.0837 0.0880
ITQ TPAMI12 0.1999 0.2821 0.2964 0.3764 0.0533 0.0765 0.0892 0.1182 0.0725 0.0899 0.0918 0.1007
IMH TIP15 0.1282 0.1536 0.1613 0.1681 0.0330 0.0361 0.0364 0.0386 0.0572 0.0668 0.0638 0.0651
PCA CG93 0.2165 0.2530 0.2701 0.2719 0.0547 0.0598 0.0632 0.0695 0.0808 0.0868 0.0891 0.0980
HashNet ICCV17 0.2086 0.2386 0.2516 0.2749 0.0528 0.0566 0.0595 0.0633 0.0746 0.0825 0.0865 0.0944
GreedyHash NeurIPS18 0.3420 0.4169 0.4240 0.4639 0.1132 0.1707 0.1841 0.2326 0.0873 0.1259 0.1398 0.1550
JMLH ICCV19 0.3607 0.4364 0.4408 0.4711 0.1078 0.1555 0.1987 0.2310 0.0860 0.1168 0.1345 0.1500
ADSH Access19 0.3360 0.4787 0.5105 0.5454 0.0858 0.1607 0.1827 0.2424 0.0886 0.1230 0.1341 0.1731
CSQ CVPR20 0.3194 0.3988 0.3773 0.4072 0.0996 0.1588 0.1712 0.2201 0.0915 0.1211 0.1375 0.1669
DPN IJCAI20 0.1783 0.2086 0.2378 0.2565 0.0445 0.0728 0.0772 0.1003 0.0737 0.0825 0.0894 0.1126
BiHalf AAAI21 0.3440 0.4036 0.4223 0.4577 0.0794 0.1280 0.1573 0.2142 0.0077 0.0490 0.0512 0.1307
OrthoCos NeurIPS21 0.1709 0.2302 0.2312 0.2566 0.0451 0.0660 0.0736 0.0984 0.0728 0.0809 0.0868 0.1126
CIBHash IJCAI21 0.2113 0.2304 0.2481 0.2618 0.0351 0.0404 0.0411 0.0453 0.0690 0.0808 0.0809 0.0883
TBH CVPR20 0.0941 0.1201 0.1073 0.1730 0.0157 0.0176 0.0226 0.0252 0.0169 0.0230 0.0248 0.0287
TSK MM16 0.2262 0.3109 0.3873 0.4151 0.0739 0.1200 0.1394 0.1112
ASZH TKDE22 0.2619 0.3787 0.4032 0.4158 0.0764 0.1192 0.1294 0.1727 - - - -
SitNet IJCAI17 0.2344 0.2406 0.2549 0.2650 0.0880 0.1127 0.1141 0.1167 - - - -
OPZH PRL19 0.1056 0.1390 0.1618 0.1961 0.0632 0.0879 0.0962 0.1143 - - - -
AH ICME17 0.2275 0.1989 0.3154 0.3557 0.0480 0.0897 0.1089 0.1445 - - - -
PIXEL CIKM24 0.3981 0.4792 0.5133 0.5465 0.1136 0.1777 0.1994 0.2519 0.0923 0.1264 0.1415 0.1741
COMAE Ours 0.3843 0.5073 0.5149 0.5801 0.1142 0.1803 0.2012 0.2532 0.0925 0.1278 0.1465 0.1764
@@ 4. Experiments
  • RQ1: Can COMAE gain superior performance in comparison to other baselines?
  • RQ2: How does COMAE perform in response to the change in the ratio of unseen classes in the training process?
  • RQ3: How does COMAE perform in case studies, where its effectiveness is evaluated in a more intuitive way?
  • RQ4: Is COMAE a running efficient hashing model?
  • RQ5: How does COMAE perform in model analysis, including the evaluation of its component change?
@@ 4.1. Experimental Settings
4.1.1. Datasets
In this paper, we analyze three distinguished zero-shot hashing datasets referenced in (Shi et al., 2022), i.e., AWA2 (Xian et al., 2018), CUB (Wah et al., 2011), SUN (Patterson and Hays, 2012). AWA2 encompasses 37,322 images across 50 animal categories, with a configuration of 40 seen and 10 unseen classes, each described by 85 attributes [39]. The CUB dataset includes 11,788 images from 200 bird species, categorized into 150 seen and 50 unseen classes, each characterized by 312 attributes. SUN dataset, on the other hand, contains 14,340 images from 717 scene categories, with a division of 645 seen and 72 unseen classes, each depicted by 102 attributes.
4.1.2. Baselines
(1) Image Shallow Hashing methods often leverage handcrafted features or employ traditional machine-learning algorithms to generate compact binary representations. Typical works include LSH (Andoni and Indyk, 2008), SH (Weiss et al., 2008), ITQ (Gong et al., 2012), IMH (Shen et al., 2015), and PCA (Maćkiewicz and Ratajczak, 1993). Shallow methods are computationally efficient and exhibit simplicity in their design, yet they may struggle to capture complex and high-level semantic features inherent in images.
(2) Image Deep Hashing methods, leverage deep neural networks to automatically learn hierarchical and abstract representations of images. Representative works can be categorized into two paths, supervised hashing works, i.e, HashNet (Cao et al., 2017), GreedyHash (Su et al., 2018), JMLH (Shen et al., 2019), ADSH (Zhou et al., 2019), CSQ (Yuan et al., 2020), DPN (Fan et al., 2020), OrthoCos (Hoe et al., 2021), and unsupervised hashing works, i.e, BiHalf (Li and van Gemert, 2021), TBH (Shen et al., 2020), CIBHash (Qiu et al., 2021). By employing end-to-end learning, deep hashing methods aim to optimize hash codes directly from raw pixel values, allowing for the discovery of intricate and discriminative image representations.
(3) Zero-shot Image Hashing methods extend the above algorithms to handle scenarios where the model encounters classes during testing that were not seen during training. These methods often incorporate semantic information, such as image attribute descriptions, or meta-learning strategies to enhance the model’s adaptability to unseen classes. Representative works include TSK (Yang et al., 2016), ASZH, SitNet, OPZH, and AH, PIXEL(Dong et al., 2024a). ASZH (Shi et al., 2022) uses category labels to learn the hash functions from the seen training data, to generate binary codes for unseen data. SitNet (Guo et al., 2017), and OPZH (Zhang et al., 2019) adopt semantic vectors to find a semantic embedding space for knowledge transferability. AH (Xu et al., 2017) constructs relationships and image features for transferability. PIXEL(Dong et al., 2024a) improves accuracy by incorporating BERT as the text encoder.
4.1.3. Evaluation Metric
To evaluate different aspects of our study, we employed a variety of metrics. These include the mean Average Precision (mAP) and the Area Under the Precision-Recall Curve (AUC) for assessing the ranking quality of search outcomes. Specifically, we present findings for the top 5,000 samples and for the entire dataset, denoted as mAP@5000 and mAP@all, respectively. Additionally, we utilized four conventional metrics: mean Average Precision (mAP), Precision-Recall Curve (PR Curve), Precision@N Curve (P@N Curve), and Recall@N Curve (R@N Curve), with a particular focus on the ranking performance for the initial 5,000 retrieved samples.
4.1.4. Implementation Details
All comparative baselines are conducted with their default hyper-parameter configurations. For a fair comparison, their image feature extractors are substituted with the ResNet101 (Khan et al., 2018). The training epoch is set to 10 and the batch size is set to 64, more than 10 will be overfitting, resulting in a decrease in accuracy. We take Adam as the optimizer weight decay 0.0005, and the learning rate is set to 0.0001. The $\epsilon$ in Eqn. 3 is set to 0.9, the $\lambda_{1},...,\lambda_{4}$ in Eqn. 3 are set to 10, 1, 10, and 1, respectively. To ensure the statistical significance of our experimental data, the average outcomes of each technique were calculated based on ten iterations.
(a) PR Curve (AWA2)
(b) PR Curve (CUB)
(c) PR Curve (SUN)
(d) P@N Curve (AWA2)
(e) P@N Curve (CUB)
(f) P@N Curve (SUN)
(g) R@N Curve (AWA2)
(h) R@N Curve (CUB)
(i) R@N Curve (SUN)
4.2. RQ1: Overall Performance
@@
Figure 3. mAP and AUC scores with the change of unseen classes ratio in the training process.
@@
Table. 1 displays the results of the COMAE and all competitive baselines on datasets AWA2, CUB, and SUN with hash code lengths of 24, 48, 64, and 128. Note that the source codes of ASZH, SitNet, AH, and OPZH are unavailable, thus we use their results in the paper (Shi et al., 2022). The experimental settings of Table 1 are the same as (Shi et al., 2022). In Fig. 2, we illustrate three distinct curves, i.e., PR Curve, R@N Curve, and the P@N Curve, to evaluate the performance of hash codes with a length of 64 bits on all datasets. Typically, the efficacy of a method is gauged by the PR curve approaching the upper right quadrant, with higher positions on the P@N and R@N curves indicative of superior performance.
@@
We find the following conclusions: (1) As shown in Table. 1, deep hashing models perform better than shallow hashing ones in general, which shows the superiority of neural networks. In addition, the zero-shot hashing models exhibit superior performance in comparison to both deep and shallow hashing models. (2) COMAE surpasses all existing zero-shot hashing benchmarks, in terms of all bit lengths. At the same time, the performance of COMAE remains relatively stable when the number of bits increases. Compared with PIXEL, the latest attribute-based zero-shot hashing method, we gain better performance, which serves to validate the correctness and effectiveness of the motivation we put forth. (3) Based on the results shown in Fig. 2 across all datasets, COMAE attains a heightened accuracy at a comparatively lower recall rate, thereby demonstrating its superiority.